OpenStudy (gabylovesyou):

What is the solution to the equation 9^(x − 2) = 27? x = 2.5 x = 0.5 x = −0.5 x = 3.5

OpenStudy (gabylovesyou):

@jim_thompson5910

satellite73 (satellite73):

use base three for both sides

satellite73 (satellite73):

since \(27=3^3\) and \(9=3^2\) you can write it as \[3^{2(x-2)}=3^3\]

OpenStudy (bmk614):

You need to take the log of both sides give you \[\log_{9}27=x-2\] And then solve for x

satellite73 (satellite73):

now that the bases are the same, set \[2(x-3)=3\] and solve for \(x\)

OpenStudy (gabylovesyou):

i get D

satellite73 (satellite73):

yeah D

satellite73 (satellite73):

mistake in the distributive law there

satellite73 (satellite73):

\[2(x-2)=3\\ 2x-4=3\] etc